Given that
ln a = 2 and ln b = 4
, use the change of base formula to find
loga(b^8).

Answer by lwsshak3(11628)

Given that
ln a = 2 and ln b = 4
use the change of base formula to find
loga(b^8).
**
loga(b^8)=ln b^8/ln a=8ln b/ln a=8*4/ln b=32/2=16